Solve. 4(7x + 10) = 3(9x – 5) + 20
Answer by checkley75(3666)

4(7x+10)=3(9x-5)+20
28x+40=27x-15+20
28x-27x=-15+20-40
x=-35 answer
proof
4(7*-35+10)=3(9*-35-5)+20
4(-245+10)=3(-315-5)+20
4(-235)=3(-320)+20
-940=-960+20
940=940